Fighter Jet pilots can be subjected to quite large forces during flight maneuvers. A 80 kg pilot is flying in a swooping arc, down and then up, with a radius of 500m and a speed of 140 m/s. What is the normal force of the seat on the pilot when she is at the lowest point of the arc?
1
Expert's answer
2020-10-29T10:17:35-0400

As per the question,

Mass of the pilot "(M)=80kg"

Radius of arc "(R)=500m"

Speed of the fighter plane "(v)= 140 m\/sec"

Let g is the gravitational acceleration, so "g=9.8 m\/sec^2"

Fighter plane is moving in the circular orbit, hence, force acting on the fighter plane will be as per the below equation.

At the lowest point, weight of the pilot is in downward direction and the pseudo of the centripetal force is also in the downward direction, hence net force applied by the pilot on the seat

"\Rightarrow F=\frac{mv^2}{R}+mg"


"\Rightarrow F =( \frac{80\times 140^2}{500}+80\times 9.8 )N"


"\Rightarrow F=(3136+784)N"


"\Rightarrow F=3920N"

Hence, normal force of seat on the pilot "(R)=F"

"(R)=3920N"